# Exact models of chiral flat-band superconductors

### v1

@@ Abstract
Recent experiments have reported evidence of chiral, time-reversal-breaking superconductivity (SC) in rhombohedral $L$-layer graphene (RLLG) with $L=4,5$ Han et al. (2025), $6$ Nguyen et al. (2025) and twisted MoTe2 (tMoTe2) Xu et al. (2025). While SC is, by now, rather common in moiré and more generally flat band (FB) systems, a remarkable feature of these superconductors is that their parent band is likely spin and valley polarized. This poses a major challenge to standard theoretical approaches to superconductivity, which typically rely on the weak-coupling instability of Fermi surfaces of time-reversal related flavors Lian et al. (2019); You and Vishwanath (2019); Yu et al. (2023); Christos et al. (2023); Putzer and Scheurer (2025); Chen et al. (2025); Schrade and Fu (2024). Together with evidence of other strongly correlated phenomena in these materials, this suggests that a strong coupling perspective, where the dispersion is considered small compared to interaction, is a better starting point. In such a FB limit, the physics is mostly determined by the interplay between the interaction and the FB wavefunctions, known as the quantum geometry (QG)  [^1] Here, we use QG to refer to the full gauge-invariant information encoded in wavefunction overlaps between arbitrary momenta rather than the long wavelength quantities such as the quantum metric.. As a result, QG is expected to control many properties of the SC phase including the pairing symmetry, single-particle gap, and phase stiffness; the latter two together limit the critical temperature of the phase. Several recent works have aimed to study the effects of QG on pairing in single-flavor bands Geier et al. (2024); Guerci et al. (2025); Wang et al. (2024a); Christos et al. (2025); Patri and Franz (2025); Shavit and Alicea (2025); Yang and Zhang (2025); May-Mann et al. (2025); Dong and Lee (2025); Crépel and Fu (2021); Guerci and Fu (2025) using various approaches. However, an analytically controlled theory beyond the weak coupling limit remains lacking.
@@
On the other hand, the development of exactly solvable models, while rare in realistic setups, can play a pivotal role in understanding complex phases of matter in FB systems. A classic example is the Haldane-Trugman-Kivelson pseudopotential model Haldane (1983); Trugman and Kivelson (1985) for the fractional quantum Hall effect in the lowest Landau level, which energetically justifies the Laughlin wavefunction Laughlin (1983) in the presence of short-ranged repulsions. This has recently been extended to a wider class of FBs, called ideal or vortexable Tarnopolsky et al. (2019); Ledwith et al. (2020); Ozawa and Mera (2021); Mera and Ozawa (2021a, b); Mera et al. ; Wang et al. (2021); Ledwith et al. (2022, 2023), allowing the development of criteria for realizing fractional Chern insulators Regnault and Bernevig (2011); Neupert et al. (2011); Sheng et al. (2011); Qi (2011); Parameswaran et al. (2012); Wu et al. (2013); Liu and Bergholtz (2022); Parameswaran et al. (2013); Bergholtz and Liu (2013) in these systems. Other examples include attractive Hubbard models with spinful time-reversal symmetry Tovmasyan et al. (2016); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022a) for FB superconductivity Peotta and Törmä (2015); Julku et al. (2016); Törmä et al. (2022); Peotta et al. (2023), where an exact $s$-wave pairing ground state is stabilized by the local attraction, as well as models for FB ferromagnetism with repulsive interactions Mielke and Tasaki (1993); Kang and Vafek (2019); Seo et al. (2019); Repellin et al. ; Bultinck et al. ; Lian et al. ; Bernevig et al. (2021); Ledwith et al. (2021); Kang et al. (2024). Recently, two of us have introduced a set of compatibility conditions of the FB wavefunctions, called ‘quantum geometric nesting’ (QGN) Han et al. (2024a) conditions (see also Zhang et al. (2025); Sun et al. (2024)), which enable systematic constructions of infinitely many solvable locally interacting models for certain symmetry broken states that unify the previous examples. One major drawback of the QGN models, however, is that the constructed order parameters are uniform in momentum space. This rules out their applicability to SC phases in a flavor-polarized FB, where the order parameter must vanish at certain points in the Brillouin zone (BZ) by the Pauli exclusion principle.
@@
where $V>0$, $\hat{n}_{\bm{R}\alpha}\equiv\hat{c}^{\dagger}_{\bm{R}\alpha}\hat{c}_{\bm{R}\alpha}$, and $\hat{H}_{AB,2}=V\sum_{\bm{R}}\left[\rho^{\text{R}}_{B}(\hat{n}_{\bm{R}A}+\hat{n}_{\bm{R}B})+\rho^{\text{FB}}_{A}\hat{n}_{\bm{R}B}\right]$ is a single-particle counter term ensuring that the projected Hamiltonian has no single-particle dispersion, i.e. it only contains fermion quartic terms $\sim\hat{\gamma}^{\dagger}\hat{\gamma}\hat{\gamma}^{\dagger}\hat{\gamma}$. $\rho^{\text{R}}_{\alpha}$ and $\rho_{\alpha}^{\text{FB}}$ are, respectively, the average electron density on $\alpha$ orbital in the occupied (lower) remote bands and the FB, e.g. $\rho_{\alpha}^{\text{FB}}\equiv\int\frac{\mathrm{d}\bm{k}}{\Omega_{\text{BZ}}}|u_{\alpha}(\bm{k})|^{2}$ where the integration is over the first BZ with volume $\Omega_{\text{BZ}}$. This counter term is a simple modification to the charge transfer gaps of the single-particle Hamiltonian and admits an apparent simplification when the FB is the lowest or the highest in the spectrum.
@@
Acknowledgement. We thank Patrick Ledwith, Ashvin Vishwanath, Qingchen Li, Tixuan Tan, Tonghang Han, Maine Christos, Zhihuan Dong and Steven A. Kivelson for the helpful discussions. Z. H. is supported by a Simons Investigator award, the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Ashvin Vishwanath, 651440). J.H.-A. is supported by a Hertz Fellowship. E. K. is supported by NSF MRSEC DMR-2308817 through the Center for Dynamics and Control of Materials. The authors thank the Harvard FAS Reaserch Computing (FASRC) for computational support.

### v2

@@ Abstract
Recent experiments have reported evidence of chiral, time-reversal-breaking superconductivity (SC) in rhombohedral $L$-layer graphene (RLLG) with $L=4,5$ Han et al. (2025), $6$ Nguyen et al. (2025) and twisted MoTe2 (tMoTe2) Xu et al. (2025). While SC is, by now, rather common in moiré and more generally flat band (FB) systems, a remarkable feature of these superconductors is that their parent band is likely spin and valley polarized. This poses a major challenge to standard theoretical approaches to superconductivity, which typically rely on the weak-coupling instability of Fermi surfaces of time-reversal related flavors Lian et al. (2019); You and Vishwanath (2019); Yu et al. (2023); Christos et al. (2023); Putzer and Scheurer (2025); Chen et al. (2025a); Schrade and Fu (2024); Jiang et al. (2025a); Chau et al. (2024). Together with evidence of other strongly correlated phenomena in these materials, this suggests that a strong coupling perspective, where the dispersion is considered small compared to interaction, is a better starting point. In such a FB limit, the physics is mostly determined by the interplay between the interaction and the FB wavefunctions, known as the quantum geometry (QG)  [^1] Here, we use QG to refer to the full gauge-invariant information encoded in wavefunction overlaps between arbitrary momenta rather than the long wavelength quantities such as the quantum metric.. As a result, QG is expected to control many properties of the SC phase including the pairing symmetry, single-particle gap, and phase stiffness; the latter two together limit the critical temperature of the phase. Several recent works have aimed to study the effects of QG on pairing in single-flavor bands Geier et al. (2024); Guerci et al. (2025); Wang et al. (2024a); Christos et al. (2025); Patri and Franz (2025); Shavit and Alicea (2025); Yang and Zhang (2025); May-Mann et al. (2025); Dong and Lee (2025); Crépel and Fu (2021); Guerci and Fu (2025); Chen et al. (2025b); Hu et al. (2025) using various approaches. However, an analytically controlled theory beyond the weak coupling limit remains lacking.
@@
On the other hand, the development of exactly solvable models, while rare in realistic setups, can play a pivotal role in understanding complex phases of matter in FB systems. A classic example is the Haldane-Trugman-Kivelson pseudopotential model Haldane (1983); Trugman and Kivelson (1985) for the fractional quantum Hall effect in the lowest Landau level, which energetically justifies the Laughlin wavefunction Laughlin (1983) in the presence of short-ranged repulsions. This has recently been extended to a wider class of FBs, called ideal or vortexable Tarnopolsky et al. (2019); Ledwith et al. (2020); Ozawa and Mera (2021); Mera and Ozawa (2021a, b); Mera et al. ; Wang et al. (2021); Ledwith et al. (2022, 2023), allowing the development of criteria for realizing fractional Chern insulators Regnault and Bernevig (2011); Neupert et al. (2011); Sheng et al. (2011); Qi (2011); Parameswaran et al. (2012); Wu et al. (2013); Liu and Bergholtz (2022); Parameswaran et al. (2013); Bergholtz and Liu (2013) in these systems. Other examples include attractive Hubbard models with spinful time-reversal symmetry Tovmasyan et al. (2016); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022a) for FB superconductivity Peotta and Törmä (2015); Julku et al. (2016); Törmä et al. (2022); Peotta et al. (2023); Jiang et al. (2025b); Jiang and Barlas (2024); Chen and Law (2024), where an exact $s$-wave pairing ground state is stabilized by the local attraction, as well as models for FB ferromagnetism with repulsive interactions Mielke and Tasaki (1993); Kang and Vafek (2019); Seo et al. (2019); Repellin et al. ; Bultinck et al. ; Lian et al. ; Bernevig et al. (2021); Ledwith et al. (2021); Kang et al. (2024). Recently, two of us have introduced a set of compatibility conditions of the FB wavefunctions, called ‘quantum geometric nesting’ (QGN) Han et al. (2024a) conditions (see also Zhang et al. (2025); Sun et al. (2024)), which enable systematic constructions of infinitely many solvable locally interacting models for certain symmetry broken states that unify the previous examples. One major drawback of the QGN models, however, is that the constructed order parameters are uniform in momentum space. This rules out their applicability to SC phases in a flavor-polarized FB, where the order parameter must vanish at certain points in the Brillouin zone (BZ) by the Pauli exclusion principle.
@@
where $V>0$, $\hat{n}_{\bm{R}\alpha}\equiv\hat{c}^{\dagger}_{\bm{R}\alpha}\hat{c}_{\bm{R}\alpha}$, and $\hat{H}_{AB,2}=V\sum_{\bm{R}}\left[\rho^{\text{R}}_{B}(\hat{n}_{\bm{R}A}+\hat{n}_{\bm{R}B})+\rho^{\text{FB}}_{A}\hat{n}_{\bm{R}B}\right]$ is a single-particle counter term ensuring that the projected Hamiltonian has no bare dispersion, i.e. it only contains fermion quartic terms $\sim\hat{\gamma}^{\dagger}\hat{\gamma}\hat{\gamma}^{\dagger}\hat{\gamma}$. $\rho^{\text{R}}_{\alpha}$ and $\rho_{\alpha}^{\text{FB}}$ are, respectively, the average electron density on $\alpha$ orbital in the occupied (lower) remote bands and the FB, e.g. $\rho_{\alpha}^{\text{FB}}\equiv\int\frac{\mathrm{d}\bm{k}}{\Omega_{\text{BZ}}}|u_{\alpha}(\bm{k})|^{2}$ where the integration is over the first BZ with volume $\Omega_{\text{BZ}}$. This counter term is a simple modification to the charge transfer gaps of the single-particle Hamiltonian and admits an apparent simplification when the FB is the lowest or the highest in the spectrum.
@@
Acknowledgement. We thank Patrick Ledwith, Ashvin Vishwanath, Qingchen Li, Tixuan Tan, Tonghang Han, Maine Christos, Zhihuan Dong and Steven A. Kivelson for the helpful discussions. We thank Felix Desrochers for pointing out a typo in the first version. Z. H. is supported by a Simons Investigator award, the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Ashvin Vishwanath, 651440). J.H.-A. is supported by a Hertz Fellowship. E. K. is supported by NSF MRSEC DMR-2308817 through the Center for Dynamics and Control of Materials. The authors thank the Harvard FAS Reaserch Computing (FASRC) for computational support.